OpenStudy (anonymous):

Whats the equation that represents the line passing through the point (1, -4) with a slope of 1/2.

hero (hero):

m(x-x1) = (y-y1)

OpenStudy (anonymous):

-4=1/2(1) + b -9/2 = b y=1/2x -9/2

OpenStudy (anonymous):

lol ima make u fail

OpenStudy (anonymous):

Wrong answer buddy.

OpenStudy (anonymous):

3.5?

OpenStudy (anonymous):

Its not one of the options

hero (hero):

y = .5x - 4.5

OpenStudy (anonymous):

haha my answer right?

hero (hero):

apparently

OpenStudy (anonymous):

Not an option either /:

hero (hero):

The answer might be in a different form

hero (hero):

mathnoob, I'm going to report you and you will be banned

hero (hero):

For saying unnecessary stuff that's disrespectful

OpenStudy (anonymous):

I agree; Hero, want me to give you the iotions?

OpenStudy (anonymous):

options*

hero (hero):

go ahead

OpenStudy (anonymous):

im mad i got it wrong

OpenStudy (anonymous):

y - 4 = 1/2(x - 1) y + 4 = 2(x - 1) y + 4 = 1/2(x - 1) y - 4 = 2(x + 1)

OpenStudy (anonymous):

wow i got banned -__-

myininaya (myininaya):

\[y-(-4)=\frac{1}{2}(x-1)\] \[y+4=\frac{1}{2}x-\frac{1}{2}\]

hero (hero):

It's the third option

OpenStudy (anonymous):

follow this equation y-y1=m(x-x1) = y-(-4)=1/2(x-1) resolve algebraically and the answer is -4.5

hero (hero):

I just simplified to y = mx+b instead of leaving it in point-slope form.

OpenStudy (anonymous):

oh i forgot in slope intercept form y=1/2x-4.5